# Energy-Based Reward Models for Robust Language Model Alignment

### v1

@@ 3.1 Reinforcement Learning from Human Feedback (RLHF)
$P(y^{+}\succ y^{-}|x)=\frac{e^{r_{\varphi}(x,y^{+})}}{e^{r_{\varphi}(x,y^{+})}+e^{r_{\varphi}(x,y^{-})}}=\sigma(r_{\varphi}(x,y^{+})-r_{\varphi}(x,y^{-})).$ (1)
@@
[…]
@@ 4.2 Energy-Based Model Formulation
A RM outputs a reward score $r=r_{\varphi}(x,y)$ that reflects how well a response $y$ to a prompt $x$ aligns with human preference. While scalar rewards are effective in some scenarios, they can be overly optimistic (see Appendix G for examples), failing to capture uncertainty and subtle distributional properties. To address this limitation post-hoc, without retraining the RM, we propose a Conditional Energy-Based Model $f_{\theta}(e,r)$, where $e=e(x,y)$ is the embedding extracted from the Base RM’s penultimate layer, and $r$ is the reward score associated with it. We define $f_{\theta}(e,r):\mathrm{R}^{d}\times\mathrm{R}\rightarrow\mathrm{R}$, where higher $f_{\theta}(e,r)$ implies greater compatibility between the embedding $e$ and the reward $r$. Our goal is to model the conditional distribution $p(r|e)$ in Eq. 4. This formulation generalizes classical models such as Bradley-Terry and Thurstone by relaxing their assumptions on the preference distribution. A detailed comparison is provided in Appendix D.
A direct maximum likelihood approach minimizes negative log-likelihood (NLL), $-\log p(r|e)$, which requires approximating the partition function $Z_{\theta}(e)$. Techniques like Importance Sampling (IS) (Gustafsson et al., 2020a) can approximate this integral. However, we found unstable or suboptimal performance in our early experiments using NLL with approximate $Z_{\theta}(e)$. It prompted us to employ Noise-Contrastive Estimation (NCE) (Gutmann & Hyvärinen, 2010), which bypasses the need to compute $Z_{\theta}(e)$ by reframing density estimation as a binary classification problem. The EBM is thus trained by maximizing the log-probability of correctly classifying real samples from “noise” samples. We generate negative samples from a Gaussian distribution centered on the observed reward score $r_{i}$:
@@
[…]
@@ 4.3 Training
The EBM’s distributional perspective allows it to “push up” energy ($-f_{\theta}$) on conflicting outputs while “pushing down” energy for compatible reward embedding pairs. By further modeling a distribution around each $r_{i}$, our EBM better accommodates the inherent noise in the Base RM’s reward predictions, enabling it to learn a more accurate reward probability distribution. We outline the full training algorithm in Algorithm 1.
@@ 4.4 Prediction
If $r_{0}$ falls within a reasonable range $[-c,c]$, we use it as the initial value in EBM: $r=r_{0}$. Otherwise, we initialize $r$ from a uniform distribution over $[-c,c]$. We set $c=2$ in this paper, selected based on the empirical reward distribution (see Table 15). This hybrid approach prevents poorly calibrated RM rewards from constraining the refinement process while still leveraging the base RM’s prior knowledge.
@@ 4.5 Why Does EBRM Improve RM Robustness and Generalization?
Figure  2(a) illustrates how EBRM progressively refines its energy landscape during training. Negative samples from $p_{N}(r|r_{i})$ force the model to distinguish between plausible and implausible rewards, sharpening the energy landscape around valid rewards while pushing away misaligned samples. At test time, Figure  2(b) illustrates a case where the Base RM assigns an incorrect ranking, but EBRM corrects it using its learned energy function. Figure  2(c) shows a case where both models are correct. This demonstrates that EBRM enhances RMs without degrading performance on correctly ranked samples.
@@ 4.6 How Does EBRM Capture Uncertainty in Preferences?
This behavior emerges from EBRM’s ability to generalize across examples. The model is not just learning from isolated data points but from the reward structure over the entire dataset. When it sees similar embeddings with diverse rewards, it infers uncertainty and learns a broader distribution. When reward signals are more consistent, it converges on narrower, more confident distributions. To gauge whether EBRM indeed learns broader distributions when preferences are ambiguous (and narrower distributions when preferences are more rigid), we computed the variance and kurtosis of reward predictions on a RewardBench subset for each task (see Appendix E for details). As illustrated in Table 1, the Chat subset shows high variance and a kurtosis value below 3, suggesting a relatively flatter distribution that reflects ambiguity in valid reward scores. Conversely, Chat Hard, Safety, and Reasoning subsets yield low variance and kurtosis values above 3, indicating more peaked distributions and thus more concentrated, confident reward estimates. A visual comparison is shown in Figure 7.
@@ 5.2 Baselines
We compare EBRM primarily against the standard RM (Base RM) and an ensemble-based approach (Coste et al., 2024). The ensemble method combines multiple reward models using: (1) Mean Optimization, averaging reward scores; (2)Worst-Case Optimization (WCO), selecting the minimum score (a conservative approach); (3) Uncertainty-Weighted Optimization (UWO), which penalizes variance among reward models. While ensemble-based approaches are not strictly post-hoc as they require multiple trained reward models, they represent a similar strategy for refining reward signals without modifying the core RM architecture. Other recent methods lack open-source code or require extensive RM retraining, making direct comparison infeasible. See Appendix E for hyperparameter selection and ablation study.
@@ 5.5 Reinforcement Learning Experiments
To assess the effectiveness of EBRM in aligning LLMs, we evaluate its impact on policy optimization. Following Coste et al. (2024), we perform 3000 steps of Proximal Policy Optimization (PPO) for alignment. Appendix B.3 provides further implementation details.
@@ Appendix C Experiments with Varying Reward Model Sizes
To further evaluate the scalability and effectiveness of EBRM, we conduct additional experiments using larger reward models on the RewardBench and RMB test sets. The goal is to assess whether EBRM maintains its improvements in robustness and generalization when applied to larger reward models. We use SFT Pythia models to obtain a 1.3B, 2.6B parameter reward models, trained following the same procedure as the 70M Base RM and on the same dataset (see Appendix B). The top layer for EBRM constitutes of about 3% of the parameter size of each Base RM. The training procedure and hyperparameters for EBRM remain consistent with our earlier experiments. The results in Table 9, 10 show that EBRM consistently improves performance, outperforming both the baseline RM and the ensemble methods in the alignment benchmarks. In Tables 11 and 12, EBRM is able to improve the performance of a 2.6B Reward Model.
We further experimented on Skywork-Reward-Llama-3.1-8B-v0.2 trained on Skywork-Reward-Preference-80K-v0.2 (Liu et al., 2024a). Results are reported in Table 13 and Table 14. Due to the wide variance in raw reward scores from this model, we normalized all reward values to the range $[-4,4]$ prior to training. The EBRM was trained for 2 epochs on this normalized data. During inference, we initialized the reward with a uniform sample from $[-1,1]$, and set the optimization hyperparameters to $\lambda=0.4$ and $\eta=0.8$. For the Reward Model Benchmark (RMB) evaluation, we used an EBRM trained for 1 epoch, with adjusted inference parameters $\lambda=0.8$ and $\eta=0.4$.
@@
Appendix D Comparison between EBRM, Bradley-Terry, and Thurstone Models
In this section, we show that the Energy-Based Reward Model (EBRM) generalizes the Bradley-Terry (Bradley & Terry, 1952) and Thurstone (Thurstone, 1927) models by relaxing their assumptions on the preference distribution.
In the Bradley-Terry model, the preference probability is modeled by assuming that the noise follows a logistic distribution:
$P(i\succ j)=\frac{e^{r_{i}}}{e^{r_{i}}+e^{r_{j}}}=\frac{1}{1+e^{-(r_{i}-r_{j})}}.$
In the Thurstone model, the reward $r_{i}$ is assumed to follow a Gaussian distribution, $r_{i}\sim\mathcal{N}(u_{i},\sigma_{i}^{2})$. Consequently, the difference $r_{i}-r_{j}$ is also Gaussian:
$r_{i}-r_{j}\sim\mathcal{N}(u_{i}-u_{j},\sigma_{i}^{2}+\sigma_{j}^{2}),$
and the corresponding preference probability is given by:
$P(i\succ j)=P(r_{i}>r_{j})=\Phi\left(\frac{u_{i}-u_{j}}{\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}}\right),$
where $\Phi$ denotes the standard Gaussian cumulative distribution function.
In contrast, EBRM does not impose a specific parametric form on the reward distribution. Instead, it models the reward using an energy-based distribution:
$p(r_{i})\propto\exp(f_{\theta}(e_{i},r_{i})),$
where $f_{\theta}$ is a neural network parameterized by $\theta$. This formulation enables EBRM to flexibly capture complex human preference distributions.
However, due to the nonparametric nature of EBRM, there is no closed-form expression for $P(i\succ j)$. It might be possible to develop approximation methods to enable training EBRM from scratch via maximum likelihood estimation. We leave this direction to future work.
Appendix E Additional Ablation Study
@@
Effect of $\sigma$ on negative distribution sampling
@@
Effect of Different $\beta$ values
@@
Table 15: Summary statistics of the Base RM reward scores on training and validation sets. The 44M Base RM reward mean and standard deviation indicates that most scores lie within $[-2.0,2.0]$, justifying our choice of initialization bounds for EBRM inference. We also include statistics for the 1.3B Base RM.
Subset Selection for reward distribution analysis
@@
Appendix F Evaluation Details

### v2

@@ 3.1 Reinforcement Learning from Human Feedback (RLHF)
$\hskip-14.22636ptP(y^{+}\succ y^{-}|x)=\frac{e^{r_{\varphi}(x,y^{+})}}{e^{r_{\varphi}(x,y^{+})}+e^{r_{\varphi}(x,y^{-})}}=\sigma(r_{\varphi}(x,y^{+})-r_{\varphi}(x,y^{-})).$ (1)
@@
$\underset{\pi}{\text{max}}\text{ }\mathbb{E}_{x\sim P,y\sim\pi(\cdot|x)}\bigg{[}r_{\varphi}(y,x)-\beta_{KL}\log\left[\frac{\pi^{\text{PPO}}(y|x)}{\pi^{\text{init}}(y|x)}\right]\bigg{]}.$ (3)
@@ 4.2 Energy-Based Model Formulation
A RM outputs a reward score $r=r_{\varphi}(x,y)$ that reflects how well a response $y$ to a prompt $x$ aligns with human preference. While scalar rewards are effective in some scenarios, they can be overly optimistic (see Appendix F for examples), failing to capture uncertainty and subtle distributional properties. To address this limitation post-hoc, without retraining the RM, we propose a Conditional Energy-Based Model $f_{\theta}(e,r)$, where $e=e(x,y)$ is the embedding extracted from the Base RM’s penultimate layer, and $r$ is the reward score associated with it. We define $f_{\theta}(e,r):\mathrm{R}^{d}\times\mathrm{R}\rightarrow\mathrm{R}$, where higher $f_{\theta}(e,r)$ implies greater compatibility between the embedding $e$ and the reward $r$. Our goal is to model the conditional distribution $p(r|e)$ in Eq. 4.
A direct maximum likelihood approach minimizes negative log-likelihood (NLL), $\log p(r|e)$, which requires approximating the partition function $Z_{\theta}(e)$. Techniques like Importance Sampling (IS) (Gustafsson et al., 2020a) can approximate this integral. However, we found unstable or suboptimal performance in our early experiments using NLL with approximate $Z_{\theta}(e)$. It prompted us to employ Noise-Contrastive Estimation (NCE) (Gutmann & Hyvärinen, 2010), which bypasses the need to compute $Z_{\theta}(e)$ by reframing density estimation as a binary classification problem. The EBM is thus trained by maximizing the log-probability of correctly classifying real samples from “noise” samples. We generate negative samples from a Gaussian distribution centered on the observed reward score $r_{i}$:
@@
$p_{\beta}(\nu)=\mathcal{N}\bigl{(}\,0,\;\beta\sigma^{2}\bigr{)}.$ (6)
@@ 4.3 Training
The EBM’s distributional perspective allows it to “push up” energy ($f_{\theta}$) on conflicting outputs while “pushing down” energy for compatible reward embedding pairs. By further modeling a distribution around each $r_{i}$, our EBM better accommodates the inherent noise in the Base RM’s reward predictions, enabling it to learn a more accurate reward probability distribution. We outline the full training algorithm in Algorithm 1.
@@ 4.4 Prediction
If $r_{0}$ falls within a reasonable range [c,c][-c,c], we use it as the initial value in EBM: $r=r_{0}$. Otherwise, we initialize $r$ from a uniform distribution over [c,c][-c,c]. We set $c=2$ in this paper, selected based on the empirical reward distribution (see Table 16). This hybrid approach prevents poorly calibrated RM rewards from constraining the refinement process while still leveraging the base RM’s prior knowledge.
@@ 4.5 Why Does EBRM Improve RM Robustness and Generalization?
Figure 2(a) illustrates how EBRM progressively refines its energy landscape during training. Negative samples from $p_{N}(r|r_{i})$ force the model to distinguish between plausible and implausible rewards, sharpening the energy landscape around valid rewards while pushing away misaligned samples. At test time, Figure 2(b) illustrates a case where the Base RM assigns an incorrect ranking, but EBRM corrects it using its learned energy function. Figure 2(c) shows a case where both models are correct. This demonstrates that EBRM enhances RMs without degrading performance on correctly ranked samples.
@@ 4.6 How Does EBRM Capture Uncertainty in Preferences?
This behavior emerges from EBRM’s ability to generalize across examples. The model is not just learning from isolated data points but from the reward structure over the entire dataset. When it sees similar embeddings with diverse rewards, it infers uncertainty and learns a broader distribution. When reward signals are more consistent, it converges on narrower, more confident distributions. To gauge whether EBRM indeed learns broader distributions when preferences are ambiguous (and narrower distributions when preferences are more rigid), we computed the variance and kurtosis of reward predictions on a RewardBench subset for each task (see Appendix D.4 for details). As illustrated in Table 1, the Chat subset shows high variance and a kurtosis value below 3, suggesting a relatively flatter distribution that reflects ambiguity in valid reward scores. Conversely, Chat Hard, Safety, and Reasoning subsets yield low variance and kurtosis values above 3, indicating more peaked distributions and thus more concentrated, confident reward estimates. A visual comparison is shown in Figure 7.
@@ 5.2 Baselines
We compare EBRM primarily against the standard RM (Base RM) and an ensemble-based approach (Coste et al., 2024). The ensemble method combines multiple reward models using: (1) Mean Optimization, averaging reward scores; (2)Worst-Case Optimization (WCO), selecting the minimum score (a conservative approach); (3) Uncertainty-Weighted Optimization (UWO), which penalizes variance among reward models. While ensemble-based approaches are not strictly post-hoc as they require multiple trained reward models, they represent a similar strategy for refining reward signals without modifying the core RM architecture. Other recent methods lack open-source code or require extensive RM retraining, making direct comparison infeasible. See Appendix D for hyperparameter selection and ablation study.
@@ 5.5 Reinforcement Learning Experiments
To assess the effectiveness of EBRM in aligning LLMs, we evaluate its impact on policy optimization. Following Coste et al. (2024), we perform 3000 steps of Proximal Policy Optimization (PPO) for alignment. Table 9 reports the average per-epoch training time for Base RM and EBRM, showing that EBRM introduces only minimal computational overhead. Appendix B.3 provides further implementation details.
@@ B.3 RL Experiment
Method Average Time
Base RM 265.60 seconds
EBRM 280.07 seconds
Table 9: Average per-epoch training time for Base RM and EBRM in the RLHF setup. EBRM adds minimal computational overhead.
@@ Appendix C Experiments with Varying Reward Model Sizes
To further evaluate the scalability and effectiveness of EBRM, we conduct additional experiments using larger reward models on the RewardBench and RMB test sets. The goal is to assess whether EBRM maintains its improvements in robustness and generalization when applied to larger reward models. We use SFT Pythia models to obtain a 1.3B, 2.6B parameter reward models, trained following the same procedure as the 70M Base RM and on the same dataset (see Appendix B). The top layer for EBRM constitutes of about 3% of the parameter size of each Base RM. The training procedure and hyperparameters for EBRM remain consistent with our earlier experiments. The results in Table 10, 11 show that EBRM consistently improves performance, outperforming both the baseline RM and the ensemble methods in the alignment benchmarks. In Tables 12 and 13, EBRM is able to improve the performance of a 2.6B Reward Model.
We further experimented on Skywork-Reward-Llama-3.1-8B-v0.2 trained on Skywork-Reward-Preference-80K-v0.2 (Liu et al., 2024a). Results are reported in Table 14 and Table 15. Due to the wide variance in raw reward scores from this model, we normalized all reward values to the range [4,4][-4,4] prior to training. The EBRM was trained for 2 epochs on this normalized data. During inference, we initialized the reward with a uniform sample from [1,1][-1,1], and set the optimization hyperparameters to $\lambda=0.4$ and $\eta=0.8$. For the Reward Model Benchmark (RMB) evaluation, we used an EBRM trained for 1 epoch, with adjusted inference parameters $\lambda=0.8$ and $\eta=0.4$.
@@
Appendix D Additional Ablation Study
@@
D.1 Effect of $\sigma$ on negative distribution sampling
@@
D.2 Effect of Different $\beta$ values
@@ D.3 Dataset Filtering
D.3.1 Rationale behind Filtering Misaligned Pairs
Filtering removes training pairs where the Base RM ranks the rejected response higher, these examples reflect conflicted or corrupted supervision either due to annotation noise or RM miscalibration. Including them can mislead learning and reinforce RM biases.
Why these pairs are not learnable: We cannot determine whether the Base RM misranked the responses, the annotation is flawed, or both. Even if an high quality external RM is used to correctly score the pairs, its score distribution would differ from the Base RM, making the signal inconsistent and uncalibrated.
Empirical comparison: As shown in Figure 6, EBRM trained on the unfiltered dataset still improves over the Base RM, but the filtered version achieves stronger alignment and more stable PPO fine-tuning, confirming that removing these corrupted pairs improves learning.
No loss of valuable signal: In the 1.3B RM setting, only 8 examples were filtered, highlighting that filtering removes conflicted and corrupted supervision.
No circular dependency: EBRM treats the Base RM outputs as noisy priors rather than fixed targets and, via noise-aware contrastive training, learns a calibrated reward distribution. This allows EBRM to correct biases rather than reinforce them.
In summary, filtering does not discard useful training signals; it removes structurally unreliable supervision, improving generalization and robustness.
@@ D.3.1 Rationale behind Filtering Misaligned Pairs
Table 16: Summary statistics of the Base RM reward scores on training and validation sets. The 44M Base RM reward mean and standard deviation indicates that most scores lie within [2.0,2.0][-2.0,2.0], justifying our choice of initialization bounds for EBRM inference. We also include statistics for the 1.3B Base RM.
D.4 Subset Selection for reward distribution analysis
@@
Appendix E Evaluation Details